2. In a piston cylinder device with 5 kg of saturated vapor at 180 °C decreases its internal energy by 1/5 as a result of con for x =32
Answer #1

Given Initial Temperature, T1 = 180°C mass of the water inside the cylinder, m = 5 kg Work tranasfer, w = 74 kJ/kg Heat trans

Quality of the steam at the end of condensation can be computed as X2 Uz - U ug - ug Substituting the known values x2 = 2066.

Substituting this Uz - U2 = 9-W - (P3V3 - P2v2) Or (uz + P303) - (uz + P2v2) = 9 - w Or hz -h2 = 9 - w Since, h = u + Pv On r

Total energy change in the system can be calculated as AE = m(uz - u) Substituting the known values AE = 5(2164 - 2582.8) = -

c)

P 2 1002.8 kPa 0.13911 0.14952 0.19384 m/kg V